OpenStudy (ballerlea):

Give the 5th and 10th term for the sequence formula below. tn = 1 + 6n

OpenStudy (tkhunny):

We might need a place to start. \(t_{1}\) maybe?

OpenStudy (anonymous):

Perhaps assume that t(1) = 1 + 6(1) = 7 and then t(2) = 1 + 6(2) = 13 t(3) = 1 + 6(3) = 19 t(4) = 1 + 6(4) = 25 t(5) = 1 + 6(5) = 31 This does not look like a typical sequence notation: e.g. t(n+1) = t(n) + 5 with t(1) given, but perhaps the question as written is correct.